OpenStudy (taylorcasey):

solve the system by elimination. 6x - y = -5 & 4x - 2y = 6 put your answer in the form [y=integer].

OpenStudy (3mar):

- Using elimination, you would to eliminate one variable from both equations, so you automatically would get one equation with one variable!

OpenStudy (mrnood):

6x-y =-5 4x -2y =6 They ask for the answer in form y= so it makes sense to eliminate the x terms. In the first equation you have 6x In the second equation you have 4x If you multiply the first equation by 2 it will have 12x If you multiply the second equation by 3 it will have 12x you can then subtract those 2 equations and 'eliminate' the x terms

OpenStudy (gluttrell):

the answer is x=-2 and y=-7.6x−y=−5;4x−2y=6 Step: Solve6x−y=−5for y: 6x−y+−6x=−5+−6x(Add -6x to both sides) −y=−6x−5 −y −1 = −6x−5 −1 (Divide both sides by -1) y=6x+5 Step: Substitute6x+5foryin4x−2y=6: 4x−2y=6 4x−2(6x+5)=6 −8x−10=6(Simplify both sides of the equation) −8x−10+10=6+10(Add 10 to both sides) −8x=16 −8x −8 = 16 −8 (Divide both sides by -8) x=−2 Step: Substitute−2forxiny=6x+5: y=6x+5 y=(6)(−2)+5 y=−7(Simplify both sides of the equation)

OpenStudy (mathmale):

@Gluttrell : Really nice explanation! But next time, please don't present the answers or the complete solution; instead, give the other person just enough guidance so that he or she can complete the problem solution himself/herself. mathmale, Moderator

OpenStudy (gluttrell):

i Know that. they can see from my example where there flaw is and how they need to fix it.

OpenStudy (mrnood):

@ Gluttrell But they dont't - all the poster will take from your answer is "y=-7" Also - your answer uses substitution method - the question states use the elimination method. My explanation is based on elimination. Also it's "their flaw" not "there flaw". Happy Holidays ;-)

ThewaffleBRO:

ello